HUBBLE SPACE TELESCOPE DAILY REPORT #2997 PERIOD COVERED: 0000Z (UTC) 11/13/01 - 0000Z (UTC) 11/14/01 Daily Status Report as of 318/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 8901 (Dark Monitor-Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. There was no anomalous activity. 1.2 Completed Five Sets of WF/PC-2 8942 (Cycle 10 Intflat Sweeps and Linearity Test) The WF/PC-2 was used to monitor the pixel-to-pixel flatfield response and provide a linearity check. These intflat sequences are done once during the year and the images will provide a backup database in the event of complete failure of the visflat lamp as well as allow monitoring of the gain ratios. There were no reported problems. 1.3 Completed Seven Sets of WF/PC-2 9042 (An Archive To Detect The Progenitors Of Massive, Core-Collapse Supernovae) The WF/PC-2 was used to search for supernovae which have massive star progenitors. The already extensive HST archive and high-resolution ground-based images of galaxies within ~20 Mpc enables us to resolve and quantify their individual bright stellar content. As massive, evolved stars are the most luminous single objects in a galaxy, the progenitors of core-collapse supernovae should be directly detectable on pre-explosion images. One Type II progenitor has been observed this year, and the investigators have proposed a short, companion WFPC2 proposal to confirm this candidate and identify a second. The observations completed nominally. 1.4 Completed Six Sets of STIS/CCD 9285 (POMS Test Proposal: STIS Non-scripted Parallel Proposal Continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. The observations were completed as planned, and no anomalies were reported. 1.5 Completed Four Sets of WF/PC-2 8936 (Cycle 10 Supplemental Darks Pt1/3) The WF/PC-2 was used to perform a dark calibration program that obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed with no reported problems. 1.6 Completed STIS/CCD 9066 (Closing in on the Hydrogen Reionization Edge of the Universe) The Space Telescope Imaging Spectrograph (CCD) was used in parallel constrain the Hydrogen reionization edge in emission that marks the transition from a neutral to a fully ionized IGM at a predicted redshifts. The proposal completed uneventfully. 1.7 Completed STIS/CCD 8611 (Observations of Nearby Type Ia Supernovae) The Space Telescope Imaging Spectrograph (CCD) was used to obtain UV spectra of five nearby {0.02 < z < 0.08} SNe Ia in the Hubble Flow. The spectra will be taken at weekly intervals over a range in time starting slightly before maximum light and extending to +30 days. These observations will accomplish the following three goals: {1} calibration of the rest frame UV light curves of SNe Ia and an assessment of their potential use as distance indicators through UV light curve shape analyses. {2} improvement in our understanding of the physics of SNe Ia, metallicity/evolutionary effects and correlations between peak brightness and UV spectral features. {3} calibration of the SNe Ia previously observed by HST at high-redshift. This data is crucial for proper cross-filter k-corrections and calibration of the supernova photometry. The observations completed with no reported anomalous activity. 1.8 Completed Two Sets of WF/PC-2 9244 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The observations completed with no anomalous activity. 1.9 Completed STIS/CCD 8903 (Bias Monitor - Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 to build up high-S/N superbiases and track the evolution of hot columns. There were no problems. 1.10 Completed WF/PC-2 9060 (Photometry of a Statistically Significant Sample of Kuiper Belt Objects) The WF/PC-2 was used to propel the physical study of KBOs forward by performing accurate photometry at V, R, and I on a sample of up to 150 KBOs. The sample is made up of objects that will be observed at thermal infrared wavelengths by SIRTF and will be used with those data to derive the first accurate diameters and albedos for a large sample of KBOs. The observations completed nominally. 1.11 Completed Two Sets of STIS/MA1/MA2 8920 (Cycle 10 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to perform the routine monitoring of the MAMA detector dark noise, and is the primary means of checking on health of the MAMA detectors systems through frequent monitoring of the background count rate. The proposal completed with no reported anomalous activity. 1.12 Completed STIS/CCD 9110 (A Search for Kuiper Belt Object Satellites) The Space Telescope Imaging Spectrograph (CCD) was used to investigate whether the large number of collisions thought to have taken place in the primordial Kuiper belt suggest that many Kuiper belt objects {KBOs} could have suffered binary-forming collisions similar to that which formed the Pluto -- Charon binary. Detection of such KBO satellites would allow measurement of KBO masses, would help to understand the past collisional environment of the Kuiper belt, and would give a context to the otherwise unique-seeming formation of the Pluto -- Charon binary. The proposal completed with no reported problems. 1.13 Completed STIS/CCD 8910 (Sparse Field CTE Test-Internal {Cycle 10}) The Space Telescope Imaging Spectrograph (CCD) was used to perform CTE measurements that are made using the "sparse field test", along both the serial and parallel axes. This program needs special commanding to provide {a} off-center MSM positionings of some slits, and {b} the ability to read out with any amplifier {A, B, C, or D}. No problems were encountered. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 10 Successful: 10 Scheduled Re-acquisitions: 4 Successful: 4 2.2 FHST Updates: Scheduled: 20 Successful: 20 2.3 Operations Notes: Using ROP SR-1A, the SSR EDAC error counter was cleared twice. HSTAR 8398 documents that SMSELEXT (MCE-A Electronics Temperature) went out-of-limits intermittently beginning at 317/164351Z, reaching a high value of 49.36 degC (limit is 49.0 degC). Per an operations note, the limit was raised to 51.5 degC. A 486 memory dump was performed at 317/2327Z and at 318/011111107Z as directed by ROP DF-01A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW